# Classification of thin Jordan schemes

### v2

@@ 3.2. Regular thin Jordan schemes
Proof of a. We have $a\diamond 1_{\Omega}=S(\omega_{0},a(1_{\Omega}(\omega_{0})))=S(\omega_{0},a(\omega_{0}))=a$. Similarly, $1_{\Omega}\diamond a=a$.
Proof of b. This follows at once from the identity $a^{t}(a(\omega_{0}))=\omega_{0}=1_{\Omega}(\omega_{0})$.
Proof of c. $a\diamond b=c\Leftrightarrow a(b(\omega_{0}))=c(\omega_{0})\Leftrightarrow b(\omega_{0})=a^{t}(c(\omega_{0}))\Leftrightarrow a^{t}\diamond c=b$.
Proof of d. We have $x(a(\omega_{0}))=b(\omega_{0})$. Therefore, $x=S(a(\omega_{0}),b(\omega_{0}))$. ∎
It follows from the above statement that $(S,\diamond)$ is a loop with the left inverse property. It is well-known that each element $a\in S$ of a loop determines two permutations $\ell_{a}$ and $r_{a}$ on $S$, defined naturally by $\ell_{a}(x)=a\diamond x$, and $r_{a}(x)=x\diamond a$. From c it follows that $\ell_{a^{t}}\ell_{a}=1_{S}$, that is, $\ell_{a^{t}}=\ell_{a}^{-1}$. It also follows from c that left division in the loop $(S,\diamond)$ is $a^{t}\diamond b$. Right division will be denoted as $b/a$, that is, $(b/a)\diamond a=b$. Later we will prove $b/a=b\diamond a^{t}$.
@@
Proof of a. If $s\neq t\in S$, then $s\cap t=\emptyset$. Thus $\sigma(s)=s(\omega_{0})\neq t(\omega_{0})=\sigma(t)$ which proves $\sigma$ is injective. To prove surjectivity, pick an arbitrary $\omega\in\Omega$ and set $s:=S(\omega_{0},\omega)$. Then $\sigma(s)=s(\omega_{0})=\omega$.
Proof of b. We must show $a\sigma=\sigma\ell_{a}$ for each $a\in S$. But this is equivalent to showing $\forall_{s\in S}:\ a(\sigma(s))=\sigma(\ell_{a}(s))$. Now $a(\sigma(s))=a(s(\omega_{0}))=(a\diamond s)(\omega_{0})=\sigma(\ell_{a}(s))$.
Proof of c. We must show $S(\sigma(u),\sigma(v))=1_{S}(v/u)=v/u$ for all $u,v\in S$. Denote $w:=v/u$. Then $w\diamond u=v$ which implies $w(u(\omega_{0}))=v(\omega_{0})$. Hence $w=S(\sigma(u),\sigma(v))$. ∎
@@
For each pair $u,v\in S$, we set $A_{u,v}:=\{w\in S\mid u\diamond(v\diamond w)=(u\diamond v)\diamond w\}$. Observe that $1_{\Omega}\in A_{u,v}$. It follows from (7) that $A_{u,v}=A_{v,u}$.
3.2.1. Properties of $\diamond$
We will now show that $(S,\diamond)$ is a loop. We start with the following assertion.
@@ 3.2.1. Properties of $\diamond$
    If $u\diamond v=v\diamond u$, then $\ell_{u}\ell_{v}=\ell_{v}\ell_{u}=\ell_{u\diamond v}$ and $A_{u,v}=S$.
@@
Proof of (a). This is an immediate consequence of (7).
Proof of (b). As $u\diamond v\neq v\diamond u$, we have $\ell_{u\diamond v}\cap\ell_{v\diamond u}=\emptyset$. But then $(u\diamond v)\diamond w\neq(v\diamond u)\diamond w$, whence the result follows once more from (7). ∎
@@
Let $u,v\in S$, and let $A:=A_{u,v}$, $\bar{A}:=S\setminus A$. Then the following hold:
@@
    $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$; in particular, $\ell_{u\diamond v}\subseteq\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}$.
@@
    The set $A$ is invariant with respect to $\ell_{u}$ and $\ell_{v}$.
@@
If $u$ and $v$ commute, then we are done by Proposition 3.7(a). So for the balance of the proof we assume $uv\neq vu$.
Proof of (a). By (6), $\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}=\ell_{u\diamond v}\cup\ell_{v\diamond u}$ while by (7), $(u\diamond v)\diamond w\in\{u\diamond(v\diamond w),v\diamond(u\diamond w)\}$. More precisely,
$(u\diamond v)\diamond w=\left\{\begin{array}[]{rl}u\diamond(v\diamond w),&w\in A,\\ v\diamond(u\diamond w),&w\not\in A\end{array}\right$.
But this is equivalent to $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}$.
Proof of (b). From (a) and the fact that $u^{t}\diamond(u\diamond v)=v$, it follows that $\ell_{v}\subseteq\ell_{u^{t}}\ell_{u\diamond v}\cup\ell_{u\diamond v}\ell_{u^{t}}$. Taking into account that $\ell_{u^{t}}=\ell_{u}^{-1}$ we now obtain
$\ell_{v}\subseteq\ell_{u}^{-1}(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\cup(\ell_{u}\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}})\ell_{u}^{-1}$
@@
$\ell_{v}\subseteq\ell_{v}1_{A}\cup\ell_{u}^{-1}\ell_{v}\ell_{u}1_{\bar{A}}\cup\ell_{u}\ell_{v}1_{A}\ell_{u}^{-1}\cup\ell_{v}\ell_{u}1_{\bar{A}}\ell_{u}^{-1}.$ (8)
@@
Together, this implies $(\ell_{v}1_{A}\cup\ell_{v}\ell_{u}1_{\bar{A}}\ell_{u}^{-1})\cap\ell_{v}=\ell_{v}$ which in turn yields $\ell_{u}1_{\bar{A}}\ell_{u}^{-1}=1_{\bar{A}}$, or equivalently, $\ell_{u}(\bar{A})=\bar{A}$. Similarly, we obtain the equality $\ell_{v}(\bar{A})=\bar{A}$ from $u\leftrightarrow v$ symmetry. ∎
It follows from Proposition 3.8(b) that $u\diamond A_{u,v}=A_{u,v}$ and $v\diamond A_{u,v}=A_{u,v}$. Together with $1_{\Omega}\in A_{u,v}$, we see that the subloop of $(S,\diamond)$ generated by $u$ and $v$ is contained in $A_{u,v}$. In particular, $u,v\in A_{u,v}$.
@@
Since $u\diamond A_{u,v}=A_{u,v}$, it follows that $u^{t}\diamond(u\diamond A_{u,v})=u^{t}\diamond A_{u,v}$. By Proposition 3.5c, $u^{t}\diamond(u\diamond A_{u,v})=A_{u,v}$ whence $u^{t}\diamond A_{u,v}=A_{u,v}$. Analogously, $v^{t}\diamond A_{u,v}=A_{u,v}$. Now since $1_{\Omega}\in A_{u,v}$, we obtain $u^{t}\diamond v^{t}\in A_{u,v}=A_{v,u}$. Therefore $(v\diamond u)\diamond(u^{t}\diamond v^{t})=v\diamond(u\diamond(u^{t}\diamond v^{t}))=v\diamond v^{t}=1_{\Omega}$. Together with $(v\diamond u)\diamond(v\diamond u)^{t}=1_{\Omega}$, we conclude that $(v\diamond u)^{t}=u^{t}\diamond v^{t}$ as desired. ∎
@@
Equality (L) follows from equation (7), while (R) and (M) follow from the inclusion $u\in A_{u,v}=A_{v,u}$. ∎
@@
Since $a^{t}=a$, each element of the loop $(S,\diamond)$ has order two. It is well known that a Moufang loop of exponent two is a group. Hence $(S,\diamond)$ is an elementary abelian $2$-group. ∎
@@ 3.2.2. Relation to alternative algebras
A loop $(S,\diamond)$ is ring-alternative (an RA-loop, for short [4]) if the loop ring $RS$ is an alternative ring for any commutative associative ring $R$. It is known that every RA-loop is Moufang but the converse is not true [4]. RA-loops were characterized in [4, Theorem 1] as follows.
@@
$\Leftarrow$” Let $(S,\diamond)$ be an RA-loop. Then $(S,\diamond)$ is a Moufang loop [4], and therefore every element $a$ has an inverse $a^{-1}$ satisfying $\forall_{a,b\in S}\ a\diamond(a^{-1}\diamond b)=b$. This implies that $\ell_{a^{-1}}=\ell_{a}^{-1}=\ell_{a}^{t}$. Together with $\ell_{1_{\Omega}}=1_{S}$, we conclude that the partition $L:=\{\ell_{a}\}_{a\in S}$ of $S^{2}$ is a rainbow. If $\diamond$ is associative, then $(L,S)$ is a thin association scheme, and therefore it is a Jordan scheme as well.
@@
$\Rightarrow$” Assume now that the permutations $L=\{\ell_{u}\}_{u\in S}$ form a Jordan scheme. Choosing $\omega_{0}=1_{\Omega}$ we obtain a binary operation $\odot$ on $L$ via
@@
By Proposition 3.11, $(L,\odot)$ is a Moufang loop. We must check the two conditions of Theorem 3.13. If $\ell_{u},\ell_{v},\ell_{w}$ associate, then they generate a subgroup222 Here we have used the fact that any subloop of a Moufang loop that is generated by three associate elements is associative. of $L$ and therefore associate in any order. Thus the first condition of Theorem 3.13 is satisfied.
@@
RA-loops were classified in [9, Theorem 1.1]. To describe them, we must first introduce a certain construction.
Let $G$ be a finite group satisfying $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ and $G^{\prime}=\{e,s\}\subseteq Z(G)$ where $e$ is the identity element of $G$. Note that according to [9], the property $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ implies $G^{\prime}=\{e,s\}\subseteq Z(G)$.
@@
$g^{*}=\left\{\begin{array}[]{cc}g&\mbox{if }g\in Z(G)\\ gs&\mbox{if }g\not\in Z(G)\end{array}\right.$.
@@
$(g,x)\diamond(h,y)=\left\{\begin{array}[]{cc}(gh,0),&x=y=0\\ (hg,1),&x=0,y=1\\ (gh^{*},1)&x=1,y=0\\ (g_{0}h^{*}g,0)&x=y=1,\end{array}\right$. (12)
@@
Theorem 3.15 ([9, Theorem 1.1]).
The algebraic structure $(L(G,*,g_{0}),\diamond)$ is a RA-loop. Moreover, any RA-loop is isomorphic to $L(G,*,g_{0})$ for suitable $G$ and $g_{0}\in Z(G)$ as described above.
@@
Proof of (a). Define $g^{*}=g\alpha(g)$ where $\alpha(g)=e$ if $g\in Z(G)$ and $\alpha(g)=s$ otherwise. Then $(gh)^{*}=h^{*}g^{*}$ is equivalent to $g,h]=\alpha(g)\alpha(h)\alpha(gh)^{-1}$.
@@
On the other hand, if $g,h$ commute then they belong to an abelian subgroup $N:=\langle g,h,Z(G)\rangle$ that is normal in $G$. As the restriction of $\alpha$ to $N$ is a homomorpism, we have $\alpha(gh)^{-1}\alpha(g)\alpha(h)=e=[g,h]$, thereby completing the proof of part (a).
Proof of (b). For any $g\in Z(G)$ and $h\in G$, it follows from the product formula (12) that $gh=hg$, $g^{*}=g$. Therefore,
$(g,0)\diamond(h,y)=\left\{\begin{array}[]{cc}gh,&y=0\\ hg,&y=1\end{array}\right.=\left\{\begin{array}[]{cc}hg,&y=0\\ hg^{*},&y=1\\ \end{array}\right.=(h,y)\diamond(g,0),$
@@
Proof of (c). Given that $L$ is an RA-loop, it is also a Moufang loop. Thus its left nucleus $N_{\lambda}(L)$ coincides with its right nucleus $N_{\rho}(L)$. We claim $N_{\rho}(L)=Z(L)$.
@@
$\mathsf{L}=\left\{\begin{array}[]{cc}(gh,0)\diamond(f,0),&x=y=0\\ (hg,1)\diamond(f,0),&x=0,y=1\\ (gh^{*},1)\diamond(f,0),&x=1,y=0\\ (g_{0}h^{*}g,0)\diamond(f,0)&x=y=1\end{array}\right.=\left\{\begin{array}[]{cc}(ghf,0),&x=y=0\\ (hgf^{*},1),&x=0,y=1\\ (gh^{*}f^{*},1),&x=1,y=0\\ (g_{0}h^{*}gf,0)&x=y=1,\end{array}\right$.
$\mathsf{R}=(g,x)\diamond\left\{\begin{array}[]{cc}(hf,0),&y=0\\ (hf^{*},1),&y=1\end{array}=\right.\left\{\begin{array}[]{cc}(ghf,0),&x=y=0\\ (hf^{*}g,1),&x=0,y=1\\ (g(hf)^{*},1)&x=1,y=0\\ (g_{0}(hf^{*})^{*}g,0)&x=y=1.\end{array}\right.$.
@@
It follows from $g_{0}\in Z(G)$ that $g_{0}^{*}=g_{0}$. Thus $\mathsf{L}=\mathsf{R}$ holds iff $fh^{*}g=gh^{*}f$. If $f\not\in Z(G)$, then choosing $h=e$ and $g\not\in C_{G}(f)$ we obtain $\mathsf{L}\neq\mathsf{R}$. If $f\in Z(G)$, then we may take $g,h\in G$ with $h^{*},g]\neq e$ to ensure $\mathsf{L}\neq\mathsf{R}$. We thereby conclude that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$. ∎

### v3

@@ 3.2. Regular thin Jordan schemes
Proof of a: We have $a\diamond 1_{\Omega}=S(\omega_{0},a(1_{\Omega}(\omega_{0})))=S(\omega_{0},a(\omega_{0}))=a$. Similarly, $1_{\Omega}\diamond a=a$.
Proof of b: This follows at once from the identity $a^{t}(a(\omega_{0}))=\omega_{0}=1_{\Omega}(\omega_{0})$.
Proof of c: $a\diamond b=c\Leftrightarrow a(b(\omega_{0}))=c(\omega_{0})\Leftrightarrow b(\omega_{0})=a^{t}(c(\omega_{0}))\Leftrightarrow a^{t}\diamond c=b$.
Proof of d: We have $x(a(\omega_{0}))=b(\omega_{0})$. Therefore, $x=S(a(\omega_{0}),b(\omega_{0}))$. ∎
We conclude from the above that $(S,\diamond)$ is a loop with the left inverse property. It is well known that each element $a\in S$ determines two permutations $\ell_{a}$ and $r_{a}$ on $S$, defined naturally by $\ell_{a}(x)=a\diamond x$ and $r_{a}(x)=x\diamond a$. From c it follows that $\ell_{a^{t}}\ell_{a}=1_{S}$, that is, $\ell_{a^{t}}=\ell_{a}^{-1}$. It also follows from c that left division in the loop $(S,\diamond)$ is $a^{t}\diamond b$. Right division will be denoted as $b/a$, that is, $(b/a)\diamond a=b$. Later we will prove $b/a=b\diamond a^{t}$.
Recall that an isotopism between two Latin squares $L:\Omega^{2}\rightarrow S$ and $L^{\prime}:{\Omega^{\prime}}^{2}\rightarrow S^{\prime}$ is a triple of bijections $\alpha\!:\!\Omega\rightarrow\Omega^{\prime}$, $\beta\!:\!\Omega\rightarrow\Omega^{\prime}$ and $\gamma\!\!:S\rightarrow S^{\prime}$ that satisfy the identity $\gamma(L(\omega_{1},\omega_{2}))=L^{\prime}(\alpha(\omega_{1}),\beta(\omega_{2}))$ for all $\omega_{1},\omega_{2}\in\Omega$.
@@
Proof of a: If $s\neq t\in S$, then $s\cap t=\emptyset$. Thus $\sigma(s)=s(\omega_{0})\neq t(\omega_{0})=\sigma(t)$ which proves $\sigma$ is injective. To prove surjectivity, pick an arbitrary $\omega\in\Omega$ and set $s:=S(\omega_{0},\omega)$. Then $\sigma(s)=s(\omega_{0})=\omega$.
Proof of b: We must show $a\sigma=\sigma\ell_{a}$ for each $a\in S$. But this is equivalent to showing $\forall_{s\in S}:\ a(\sigma(s))=\sigma(\ell_{a}(s))$. Now $a(\sigma(s))=a(s(\omega_{0}))=(a\diamond s)(\omega_{0})=\sigma(\ell_{a}(s))$.
Proof of c: We must show $S(\sigma(u),\sigma(v))=1_{S}(v/u)=v/u$ for all $u,v\in S$. Denote $w:=v/u$. Then $w\diamond u=v$ which implies $w(u(\omega_{0}))=v(\omega_{0})$. Hence $w=S(\sigma(u),\sigma(v))$. ∎
@@
For each pair $u,v\in S$, we set ${\mathfrak{a}}({u,v}):=\{w\in S\mid u\diamond(v\diamond w)=(u\diamond v)\diamond w\}$. Observe that $1_{\Omega}\in\mathfrak{a}({u,v})$. It follows from (7) that $\mathfrak{a}({u,v})=\mathfrak{a}({v,u})$.
3.2.1. $(S,\diamond)$ is a Moufang loop
To show this we start with the following assertion:
@@ 3.2.1. $(S,\diamond)$ is a Moufang loop
    If $u\diamond v=v\diamond u$, then $\ell_{u}\ell_{v}=\ell_{v}\ell_{u}=\ell_{u\diamond v}$ and $\mathfrak{a}({u,v})=S$.
@@
Proof of a: This is an immediate consequence of (7).
Proof of b: As $u\diamond v\neq v\diamond u$, we have $\ell_{u\diamond v}\cap\ell_{v\diamond u}=\emptyset$. But then $(u\diamond v)\diamond w\neq(v\diamond u)\diamond w$ holds for every $w\in S$, whence the result follows once more from (7). ∎
@@
Let $u,v\in S$, and let $\mathfrak{a}:=\mathfrak{a}({u,v})$ and $\overline{\mathfrak{a}}:=S\setminus\mathfrak{a}$. Then the following hold:
@@
    $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{\mathfrak{a}}\cup\ell_{v}\ell_{u}1_{\overline{\mathfrak{a}}}$; in particular, $\ell_{u\diamond v}\subseteq\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}$.
@@
    The set $\mathfrak{a}$ is invariant with respect to $\ell_{u}$ and $\ell_{v}$.
@@
If $u$ and $v$ commute, then we are done by Proposition 3.8(a). So for the balance of the proof we assume $u\diamond v\neq v\diamond u$.
Proof of a: By (6), $\ell_{u}\ell_{v}\cup\ell_{v}\ell_{u}=\ell_{u\diamond v}\cup\ell_{v\diamond u}$ while by (7), $(u\diamond v)\diamond w\in\{u\diamond(v\diamond w),v\diamond(u\diamond w)\}$. More precisely,
$(u\diamond v)\diamond w=\left\{\begin{array}[]{rl}u\diamond(v\diamond w),&w\in\mathfrak{a},\\ v\diamond(u\diamond w),&w\in\overline{\mathfrak{a}}.\end{array}\right$.
But this is equivalent to $\ell_{u\diamond v}=\ell_{u}\ell_{v}1_{\mathfrak{a}}\cup\ell_{v}\ell_{u}1_{\overline{\mathfrak{a}}}$.
Proof of b: From (a) and the fact that $u^{t}\diamond(u\diamond v)=v$, it follows that $\ell_{v}\subseteq\ell_{u^{t}}\ell_{u\diamond v}\cup\ell_{u\diamond v}\ell_{u^{t}}$. Taking into account that $\ell_{u^{t}}=\ell_{u}^{-1}$ we now obtain
$\ell_{v}\subseteq\ell_{u}^{-1}(\ell_{u}\ell_{v}1_{\mathfrak{a}}\cup\ell_{v}\ell_{u}1_{\overline{\mathfrak{a}}})\cup(\ell_{u}\ell_{v}1_{\mathfrak{a}}\cup\ell_{v}\ell_{u}1_{\overline{\mathfrak{a}}})\ell_{u}^{-1}$
@@
$\ell_{v}\subseteq\ell_{v}1_{\mathfrak{a}}\cup\ell_{u}^{-1}\ell_{v}\ell_{u}1_{\overline{\mathfrak{a}}}\cup\ell_{u}\ell_{v}1_{\mathfrak{a}}\ell_{u}^{-1}\cup\ell_{v}\ell_{u}1_{\overline{\mathfrak{a}}}\ell_{u}^{-1}.$ (8)
@@
Together, this implies $(\ell_{v}1_{\mathfrak{a}}\cup\ell_{v}\ell_{u}1_{\overline{\mathfrak{a}}}\ell_{u}^{-1})\cap\ell_{v}=\ell_{v}$ which in turn yields $\ell_{u}1_{\overline{\mathfrak{a}}}\ell_{u}^{-1}=1_{\overline{\mathfrak{a}}}$, or equivalently, $\ell_{u}(\overline{\mathfrak{a}})=\overline{\mathfrak{a}}$. Similarly, we obtain the equality $\ell_{v}(\overline{\mathfrak{a}})=\overline{\mathfrak{a}}$ from $u\leftrightarrow v$ symmetry. ∎
It follows from Proposition 3.9(b) that $u\diamond\mathfrak{a}({u,v})=\mathfrak{a}({u,v})$ and $v\diamond\mathfrak{a}({u,v})=\mathfrak{a}({u,v})$. Together with $1_{\Omega}\in\mathfrak{a}({u,v})$, we see that the subloop of $(S,\diamond)$ generated by $u$ and $v$ is contained in $\mathfrak{a}({u,v})$. In particular, $u,v\in\mathfrak{a}({u,v})$.
@@
Since $u\diamond\mathfrak{a}({u,v})=\mathfrak{a}({u,v})$, it follows that $u^{t}\diamond(u\diamond\mathfrak{a}({u,v}))=u^{t}\diamond\mathfrak{a}({u,v})$. By Proposition 3.6c, $u^{t}\diamond(u\diamond\mathfrak{a}({u,v}))=\mathfrak{a}({u,v})$ whence $u^{t}\diamond\mathfrak{a}(u,v)=\mathfrak{a}({u,v})$. Analogously, $v^{t}\diamond\mathfrak{a}(u,v)=\mathfrak{a}({u,v})$. Now since $1_{\Omega}\in\mathfrak{a}(u,v)$, we obtain $u^{t}\diamond v^{t}\in\mathfrak{a}({u,v})=\mathfrak{a}({v,u})$. Therefore $(v\diamond u)\diamond(u^{t}\diamond v^{t})=v\diamond(u\diamond(u^{t}\diamond v^{t}))=v\diamond v^{t}=1_{\Omega}$. Together with $(v\diamond u)\diamond(v\diamond u)^{t}=1_{\Omega}$, we conclude that $(v\diamond u)^{t}=u^{t}\diamond v^{t}$ as desired. ∎
@@
Equality (L) follows from equation (7), while (R) and (M) follow from the inclusion $u\in\mathfrak{a}({u,v})=\mathfrak{a}({v,u})$. ∎
@@
Since $a^{t}=a$, each element of the loop $(S,\diamond)$ has order two. It is well known that a Moufang loop of exponent two is a group (cf. [5, Proposition 2]). Hence $(S,\diamond)$ is an elementary abelian $2$-group. ∎
@@ 3.2.2. Relation to alternative algebras
A loop $(S,\diamond)$ is ring-alternative (an RA-loop, for short [6]) if the loop ring $RS$ is an alternative ring for any commutative associative ring $R$. It is known that every RA-loop is Moufang but the converse is not true [6]. RA-loops were characterized in [6, Theorem 1] as follows:
@@
($\Leftarrow$) Let $(S,\diamond)$ be an RA-loop. Then $(S,\diamond)$ is a Moufang loop [6], and therefore every element $a$ has an inverse $a^{-1}$ satisfying $\forall_{a,b\in S}\ a\diamond(a^{-1}\diamond b)=b$. This implies that $\ell_{a^{-1}}=\ell_{a}^{-1}=\ell_{a}^{\,t}$. Together with $\ell_{1_{\Omega}}=1_{S}$, we conclude that the partition $L:=\{\ell_{a}\}_{a\in S}$ of $S^{2}$ is a rainbow. If $\diamond$ is associative, then $(L,S)$ is a thin association scheme, and therefore it is a Jordan scheme as well.
@@
($\Rightarrow$) Assume now that the permutations $L=\{\ell_{u}\}_{u\in S}$ form a Jordan scheme. Choosing $\omega_{0}=1_{\Omega}$ we obtain a binary operation $\odot$ on $L$ via
@@
By Proposition 3.12, $(L,\odot)$ is a Moufang loop. We must check the two conditions of Theorem 3.14. If $\ell_{u},\ell_{v},\ell_{w}$ associate, then they generate a subgroup of $L$ and therefore associate in any order. (Here we have used the fact that any subloop of a Moufang loop that is generated by three associate elements is associative.) Thus the first condition of Theorem 3.14 is satisfied.
@@
We end this subsection with an isomorphism criterion for loop-based Jordan schemes. Following [9], a half-isomorphism between two loops $(L,\diamond)$ and $(L^{\prime},\diamond^{\prime})$ is a bijection $\varphi:L\rightarrow L^{\prime}$ satisfying $\varphi(x\diamond y)\in\{\varphi(x)\diamond^{\prime}\varphi(y),\varphi(y)\diamond^{\prime}\varphi(x)\}$. Note that every loop isomorphism or anti-isomorphism is a half-isomorphism, although we regard these as trivial.
Proposition 3.16.
Given two RA-loops $(L,\diamond)$ and $(L^{\prime},\diamond^{\prime})$, their corresponding Jordan schemes $\mathcal{J}:=(L,\{\ell_{a}\}_{a\in L})$ and $\mathcal{J}^{\prime}:=(L^{\prime},\{\ell^{\prime}_{a}\}_{a\in L^{\prime}})$ are isomorphic (as Jordan schemes) if and only if $L$ and $L^{\prime}$ are half-isomorphic.
Proof.
The reverse direction is obvious. To prove the forward direction, we assume that $\mathcal{J}$ and $\mathcal{J}^{\prime}$ are isomorphic Jordan schemes (i.e. rainbows). In other words, there exists a bijection $f:L\rightarrow L^{\prime}$ such that $f\{\ell_{a}\}_{a\in L}f^{-1}=\{\ell^{\prime}_{a}\}_{a\in L^{\prime}}$. The latter is equivalent to the existence of a bijection $\varphi:L\rightarrow L^{\prime}$ satisfying $f\ell_{a}f^{-1}=\ell^{\prime}_{\varphi(a)}$ for all $a\in L$. This yields
$f(\ell_{a}\ell_{b}\cup\ell_{b}\ell_{a})f^{-1}=\ell^{\prime}_{\varphi(a)}\ell^{\prime}_{\varphi(b)}\cup\ell^{\prime}_{\varphi(b)}\ell^{\prime}_{\varphi(a)}.$
Since both $\mathcal{J}$ and $\mathcal{J}^{\prime}$ are Jordan schemes, the above equality is equivalent to
$f(\ell_{a\diamond b}\cup\ell_{b\diamond a})f^{-1}=\ell^{\prime}_{\varphi(a)\diamond^{\prime}\varphi(b)}\cup\ell^{\prime}_{\varphi(b)\diamond^{\prime}\varphi(a)}.$
Since the left-hand side of this equality is equal to $\ell^{\prime}_{\varphi(a\diamond b)}\cup\ell^{\prime}_{\varphi(b\diamond a)}$, we conclude that
$\{\varphi(a\diamond b),\varphi(b\diamond a)\}=\{\varphi(a)\diamond^{\prime}\varphi(b),\varphi(b)\diamond^{\prime}\varphi(a)\},$
as desired. ∎
Although there exist Moufang loops that are non-trivially half-isomorphic, we are not aware of any examples of isomorphic Jordan schemes based on nontrivial half-isomorphic Moufang loops.
@@
RA-loops were classified in [12, Theorem 1.1]. To describe them, we must first introduce a special construction.
Let $G$ be a finite group satisfying $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ and $G^{\prime}=\{e,s\}\subseteq Z(G)$ where $e$ is the identity element of $G$. Note that according to [12], the property $G/Z(G)\cong{\mathbb{Z}}_{2}^{2}$ implies $G^{\prime}=\{e,s\}\subseteq Z(G)$. Note also that in this case there exists a $2$-subgroup $P$ of $G$ and an abelian subgroup $A$ of $G$ such that $G=P\times A$ and $P/Z(P)\cong{\mathbb{Z}}_{2}^{2}$.
@@
$g^{*}=\left\{\begin{array}[]{ll}g,&g\in Z(G)\\ gs,&g\not\in Z(G).\end{array}\right$.
@@
$(g,x)\diamond(h,y)=\left\{\begin{array}[]{ll}(gh,0),&x=y=0\\ (hg,1),&x=0,y=1\\ (gh^{*},1),&x=1,y=0\\ (g_{0}h^{*}g,0),&x=y=1.\end{array}\right$. (12)
@@
Theorem 3.17 ([12, Theorem 1.1]).
The algebraic structure $(L(G,*,g_{0}),\diamond)$ is an RA-loop. Moreover, any RA-loop is isomorphic to $L(G,*,g_{0})$ for suitable $G$ and $g_{0}\in Z(G)$ as described above.
@@
Proof of a: Define $g^{*}=g\alpha(g)$ where $\alpha(g)=e$ if $g\in Z(G)$ and $\alpha(g)=s$ otherwise. We must show $(gh)^{*}=h^{*}g^{*}$, which is equivalent to showing $g,h]=\alpha(g)\alpha(h)\alpha(gh)^{-1}$.
@@
On the other hand, if $g,h$ commute then they belong to an abelian subgroup $N:=\langle g,h,Z(G)\rangle$ that is normal in $G$. As the restriction of $\alpha$ to $N$ is a homomorphism, we have $\alpha(gh)^{-1}\alpha(g)\alpha(h)=e=[g,h]$, thereby completing the proof of part (a).
Proof of b: For any $g\in Z(G)$ and $h\in G$, it follows from the product formula (12) that $gh=hg$, $g^{*}=g$. Therefore,
$(g,0)\diamond(h,y)=\left\{\begin{array}[]{ll}gh,&y=0\\ hg,&y=1\end{array}\right.=\left\{\begin{array}[]{ll}hg,&y=0\\ hg^{*},&y=1\\ \end{array}\right.=(h,y)\diamond(g,0),$
@@
Proof of c: Given that $L$ is an RA-loop, it is also a Moufang loop. Thus its left nucleus $N_{\lambda}(L)$ coincides with its right nucleus $N_{\rho}(L)$. We claim $N_{\rho}(L)=Z(L)$.
@@
$\mathsf{L}=\left\{\begin{array}[]{ll}(gh,0)\diamond(f,0),&x=y=0\\ (hg,1)\diamond(f,0),&x=0,y=1\\ (gh^{*},1)\diamond(f,0),&x=1,y=0\\ (g_{0}h^{*}g,0)\diamond(f,0),&x=y=1\end{array}\right.=\left\{\begin{array}[]{ll}(ghf,0),&x=y=0\\ (hgf^{*},1),&x=0,y=1\\ (gh^{*}f^{*},1),&x=1,y=0\\ (g_{0}h^{*}gf,0),&x=y=1.\end{array}\right$.
$\mathsf{R}=(g,x)\diamond\left\{\begin{array}[]{ll}(hf,0),&y=0\\ (hf^{*},1),&y=1\end{array}=\right.\left\{\begin{array}[]{ll}(ghf,0),&x=y=0\\ (hf^{*}g,1),&x=0,y=1\\ (g(hf)^{*},1),&x=1,y=0\\ (g_{0}(hf^{*})^{*}g,0),&x=y=1.\end{array}\right$.
@@
It follows from $g_{0}\in Z(G)$ that $g_{0}^{*}=g_{0}$. Thus $\mathsf{L}=\mathsf{R}$ holds iff $fh^{*}g=gh^{*}f$. If $f\not\in Z(G)$, then choosing $h=e$ and $g\not\in C_{G}(f)$ we obtain $\mathsf{L}\neq\mathsf{R}$. (Here $C_{G}(f)$ denotes the centralizer in $G$ of $f$.) If $f\in Z(G)$, then we may take $g,h\in G$ with $h^{*},g]\neq e$ to ensure $\mathsf{L}\neq\mathsf{R}$. We thereby conclude that $(f,1)\not\in N_{\rho}(L)$ for any $f\in G$. ∎